OpenStudy (anonymous):

Which of the following is an example of the associative property? A. 2(5 - 3) = 10 - 6 B. -7 + (19 + 5) = (-7 + 19) + 5 C.11 + (6 + 8) = 11 + (28 + 6) D. 1 over 3• 3 = 1 PLEASE HELP ME.

OpenStudy (anonymous):

tthe answer is b

OpenStudy (anonymous):

I agree...it is B

OpenStudy (anonymous):

it would equal B. -7 + (19 + 5) = (-7 + 19) + 5

OpenStudy (calculusxy):

The answer is B because the associative property is just the movement of the parentheses.

OpenStudy (anonymous):

calculusxy was right. In associative property only the parentheses change, the given values should not change.By looking this example you can understand (a+b)+c=a+(b+c) .

OpenStudy (anonymous):

it's b